# Uncovering Hidden Connections: Iterative Tracking and Reasoning for Video-grounded Dialog

### v1

@@ Abstract
In contrast to conventional visual question answering, video-grounded dialog necessitates a profound understanding of both dialog history and video content for accurate response generation. Despite commendable strides made by existing methodologies, they often grapple with the challenges of incrementally understanding intricate dialog histories and assimilating video information. In response to this gap, we present an iterative tracking and reasoning strategy that amalgamates a textual encoder, a visual encoder, and a generator. At its core, our textual encoder is fortified with a path tracking and aggregation mechanism, adept at gleaning nuances from dialog history that are pivotal to deciphering the posed questions. Concurrently, our visual encoder harnesses an iterative reasoning network, meticulously crafted to distill and emphasize critical visual markers from videos, enhancing the depth of visual comprehension. Culminating this enriched information, we employ the pre-trained GPT-2 model as our response generator, stitching together coherent and contextually apt answers. Our empirical assessments, conducted on two renowned datasets, testify to the prowess and adaptability of our proposed design.
@@
Video-grounded Dialog, Path Tracking, Iterative Reasoning.
@@ 1 Introduction
With the rapid proliferation of multimedia data, research attention has intensified on visual-language tasks, attesting to their pivotal role in academic research [1, 2, 3]. Notable advancements have been realized in fields such as Visual Question Answering (VQA) [4, 5], Visual Dialog [6, 7], and Video Captioning [8, 9]. Central to these tasks is the aim to refine machine capabilities in understanding visual content and to convey these understandings linguistically in ways intelligible to humans.
@@
An emerging domain of significance is video-grounded dialog [^1] https://video-dialog.com/, sometimes referred to as audio-visual scene-aware dialog by certain academicians [10, 11, 12, 13]. As illustrated in Figure 1, this rapidly evolving domain mandates the creation of an agent adept at conducting coherent dialogs rooted in video data, which is a monumental leap toward genuine human-computer synergy. The ramifications of video-grounded dialog are broad, spanning enhancements in virtual personal assistants to innovations tailored for visually impaired individuals, thereby underlining its academic importance. Spurred by advancements in deep learning, the domain of video-grounded dialog is witnessing considerable traction. Existing techniques within this sphere can be methodically categorized into:
@@
1) Textual Information Modeling. This branch emphasizes modeling textual elements, predominantly involving dialog inquiries and their preceding contexts [14, 15, 16]. For instance, Le et al. [15] devised a semantic graph anchored in the lexical components of Q&A pairs, leading to a model proficient in forecasting reasoning pathways on the graph. This methodology is adept at extracting question-specific data from dialogic contexts. In a related development, Kim et al. [14] employed the Gumbel-Softmax technique to identify dialog history utterances that are most pertinent to a given question, enhancing question representations in the process.
@@
2) Visual Information Modeling. This avenue of inquiry highlights visual elements, primarily focusing on video frames and notable objects [17, 18, 19, 20]. As a case in point, Le et al. [18] propounded a bidirectional model that encapsulates both spatial-to-temporal relations and their converse, aiming to harness the intricate spatial and temporal subtleties embedded in video. Concurrently, Geng et al. [17] advocated representing videos via spatio-temporal scene graphs, capturing salient audio-visual cues and their semantic interrelations effectively.
@@
3) Decoder Generation. This frontier emphasizes the formulation of answer derivations [21, 22, 23]. Noteworthy is Le et al. [22] extending the established functionalities of the GPT-2 architecture, eschewing conventional auto-regressive decoding strategies. Complementing this domain, the integration of pointer networks [21] has amplified the generative prowess of dialog systems.
@@
Despite the significant progress observed in the domain of video-grounded dialog, there remain pressing challenges warranting meticulous consideration: 1) Dialog History Understanding. Notable recent studies [16, 15] predominantly focus on dialog history segments that align directly with user queries. Such a concentrated perspective can inadvertently exclude broader dialog contexts, leading to the omission of significant information. The inherent lack of transparency in these methods further impedes a clear understanding of the decision-making process of the model. The imperative, therefore, is to develop a superior textual encoder, adept at a comprehensive examination of dialog histories while enhancing model interpretability. 2) Visual Content Understanding. While several works [17] endeavor to encapsulate visual signals through complex graph structures, the critical function of user-posed questions in steering visual comprehension is frequently disregarded. This omission leads to a failure to capture question-specific insights, thereby limiting the capacity to mine meaningful visual information. Additionally, the current trend of depending solely on a singular extraction process [19, 20] may be inadequate given the multi-faceted nature of video data. Therefore, an approach that harmonizes user queries with iterative reasoning is crucial for a nuanced comprehension of visual content.
@@
In response to the outlined challenges, we introduce the Iterative Tracking and Reasoning (ITR) framework for video-grounded dialog, detailed in Figure 2. This novel approach aims to redress the limitations of current methodologies by integrating an explicit history modeling mechanism alongside an iterative reasoning strategy. To be specific, we adopt a transparent path tracking and aggregation mechanism, adept at assembling salient information germane to user queries. Such a system amplifies query understanding by sourcing pivotal contextual elements from the dialog history. Concurrently, to harness the latent visual semantics in videos, a multimodal iterative reasoning network is devised. This architecture enables robust iterative reasoning processes, progressively refining the apprehension of visual content and consequently bolstering generation prowess. The prowess of the ITR paradigm is corroborated through rigorous empirical evaluations on two benchmark datasets. Experimental findings accentuate the proficiency of our framework in obviating the limitations intrinsic to prevailing strategies, thereby achieving state-of-the-art performance. To advance collaborative research, the codebase for the ITR framework is now available at https://github.com/Hyu-Zhang/ITR.
@@
  • We design an explicit path tracking and aggregation system. This methodology enables the effective extraction of vital information from previous dialog interactions, thereby fostering a nuanced understanding of user queries.
  • We devise a multimodal iterative reasoning network that leverages user questions. This feature serves to incrementally refine the comprehension of visual content, leading to more precise and context-aware dialog generation.
@@ 2.1 Visual Dialog
Visual dialog represents a sophisticated extension of the VQA framework, requiring multiple rounds of sequential interactions. The fundamental goal of this paradigm is to facilitate a flawless amalgamation of cross-modal information derived from various inputs, culminating in the formation of a coherent textual response. Current methodologies operating within this sphere may be broadly grouped into three distinct paradigms: attention mechanism-based strategies, graph network-based techniques, and approaches reliant on pre-trained models.
@@
Attention Mechanism-based Approaches. These methodologies deploy attention mechanisms to pinpoint the visual regions that bear the highest relevance to the posed question. The underlying process requires an acute understanding of the immediate context of the inquiry, as well as an insight into the historical dialog trajectory. A salient example of this approach is the Recursive visual Attention (RvA) model, formulated by Niu et al. [24]. Through iterative refinement of visual attention and retrospective analysis of dialog sequences, this model adeptly addresses co-referential ambiguities. Further complementing this area, Yang et al. [25] have innovated two Sequential Dialog Networks (SeqDialN). These networks artfully separate the complexity inherent in multimodal feature fusion from the inference process, thereby simplifying the design and implementation of the inference engine.
@@
Graph Network-based Techniques. Recognizing the pivotal role of seamless information integration, the academic community is incrementally gravitating toward graph network methodologies [26]. For instance, Schwartz et al. [27] brought forth the Factor Graph Attention (FGA) architecture. This methodology enables the meticulous construction of graphs atop utility representations, subsequently discerning their interconnections. In a parallel vein, Guo et al. [28] introduced the Context-Aware Graph (CAG) neural structure, characterized by the dynamic interdependencies exhibited by individual nodes within the graph. Uniquely, this paradigm ensures that only nodes bearing the utmost relevance are integrated, yielding a refined, context-sensitive relational graph logic.
@@
Pre-trained Model-dependent Methods. The contemporary epoch of machine learning has been marked by the ascent of pre-trained models, lauded for their adeptness in seamlessly integrating diverse informational modalities. Anchoring on this insight, Wang et al. [29] delineated a holistic strategy that harnesses pre-trained architectures to adroitly weave together visual and dialogual elements. Mirroring this trajectory, Nguyen et al. [30] unveiled the Light-weight Transformer for Many Inputs (LTMI), a pioneering neural construct that leans on an agile transformer mechanism to masterfully orchestrate interactions across multifarious utilities.
@@
However, the nuances and complexities of video-grounded dialog tasks preclude the direct application of methodologies cultivated for image-based visual dialog. Video-grounded dialog manifests longer utterance sentences, additional dialog rounds, and a more intricate context that demands sophisticated comprehension [31]. These attributes compound the challenge, rendering the direct transposition of existing visual dialog techniques impractical and inefficacious.
@@ 2.2 Video-grounded Dialog
In recent years, the academic community has observed an escalating focus on video-grounded dialog systems. Noteworthy challenges, including DSTC7 [31] and DSTC8 [32], have been instrumental in formulating benchmarks in this domain. This burgeoning field has subsequently witnessed the advent of numerous sophisticated methodologies to address its inherent complexities. For example, Hori et al. [10] pioneered an approach employing an LSTM-driven encoder-decoder framework, placing emphasis on multimodal attention to harmoniously integrate textual and visual cues. Expanding on this work, Le et al. [11] presented the Multimodal Transformer Network (MTN), devised specifically to encode visual content while concurrently integrating diverse data modalities. Similarly, innovative contributions like the Bi-directional Spatio-Temporal reasoning model (BiST)[18] have made strides in extracting both visual and spatial cues from videos, thus elucidating the interconnectedness of text and visual sequences in multidimensional aspects. In pursuit of refining co-referential understanding, Kim et al. [14] introduced the Structured Co-reference Graph Attention model (SCGA), laying the foundation for a graph-based approach deeply rooted in multimodal co-referencing techniques. In tandem with this, Geng et al. [17] advanced the Spatio-Temporal Scene Graph Representation learning technique (STSGR)—an innovative proposition capturing the semantic crux of videos through scene graphs, all the while maintaining efficient encoding.
@@
The ascendance of pre-trained language models in myriad natural language processing domains is unmistakable. Recent efforts are directed towards fusing these models with video-grounded dialog systems. To this end, Le et al. [22] advanced the VGD-GPT framework, leveraging the capabilities of the well-regarded GPT-2 model, thus transforming video-grounded dialog generation into a more fluid sequence-to-sequence task. Building on this progression, Li et al. [12] incorporated a multi-task learning approach with pre-existing language models, culminating in the inception of the Response Language Model (RLM). Recognizing the hurdles of text hallucination in text generation, Yoon et al. [23] crafted the Text HallucinAtion Mitigating (THAM) architecture, integrating Text Hallucination Regularization (THR) loss into expansive pre-trained language models. In a stride towards achieving nuanced dialog generation, Zhao et al. [33] advocated for a Multi-Agent Reinforcement Learning (MARL) methodology, accentuating sophisticated collaborations in cross-modal analysis. With the advent of Large Language Models, such as Video-LLaMA [34] and Otter [35], it is evident that this intricate domain continues to be at the forefront of research attention.
@@
To sum up, within the domain of history modeling, a prevailing trend involves the deployment of intricate graph networks to discern semantic nuances intertwined within historical data. However, these methods often compromise on interpretability. In contrast, our approach embraces a streamlined path tracking and aggregation mechanism, offering superior interpretability. Additionally, current methodologies largely overlook the iterative interplay between videos and questions. We address this oversight, devising an innovative iterative reasoning network that prioritizes this symbiotic relationship.
@@ 3.1 Problem Formulation
The task of video-grounded dialog centers on generating a relevant response $A_{t}$ in answer to a given question $Q_{t}$. This response is derived through a thorough analysis of a video $V$, a caption $C$ crafted by a human to describe the video, and its associated dialog history $H$. The dialog history $H$ comprises a series of prior questions and their answers, denoted as $(Q_{i},A_{i})$ for each $i$ ranging from $1$ to $t-1$. Specifically, $H$ is articulated as $\{(Q_{1},A_{1}),...,(Q_{t-1},A_{t-1})\}$. The holistic consideration of these elements is crucial to successfully address the posed question.
@@ 3.2.1 Preprocessing
The word embeddings for the given question $Q_{t}$, the caption $C$, and the dialog history $H$ are procured via a consistent mechanism. To elucidate, let’s consider the question $Q_{t}=\{w_{1},w_{2},...,w_{n_{q}}\}$ as an example. Here, an operation $f$ is applied to the vocabulary obtained from the word-to-id mapping function, Vocab($\cdot$). This process is represented by the following expression:
@@
where $\mathbf{E}_{q}\in\mathbb{R}^{{n_{q}}\times{d}}$ signifies the resulting word embeddings of the question. The function $f$ corresponds to retrieving the embedding vector associated with the id from the pre-established embedding matrix. The same technique is applied to ascertain the feature of the caption and the dialog history, symbolized as $\mathbf{E}_{c}\in\mathbb{R}^{n_{c}\times d}$ and $\mathbf{E}_{h}\in\mathbb{R}^{(t-1)\times n_{h}\times d}$, wherein $n_{c}$ and $n_{h}$ respectively designates the number of words present in the caption and each utterance.
@@
To bolster the interpretability of questions drawn from the dialog history [^2] Given that the caption inherently encapsulates the essential information, it is utilized directly in the decoding generation process without necessitating further operations., we introduce two preprocessing steps: co-reference resolution and dependency parsing. Initially, the dialog history is integrated with the present question to create a unified sequence. We then utilize the AllenNLP [^3] https://allenai.org/allennlp. toolkit for co-reference resolution, aiming to dispel potential confusion related to pronouns. Following this, the Stanford CoreNLP utility [^4] https://stanfordnlp.github.io/CoreNLP/. is employed to derive the dependency parse tree. Our attention is concentrated on the dialog history responses due to the inherent similarities between questions and their corresponding answers. This tree delineates the syntactic relationships between sentence constituents. By harnessing the subject-verb-object paradigm, we craft pertinent triplets in the $\langle{subject,relation,object}\rangle$ format. This conversion facilitates the representation of both the question and dialog history as sets of triplet pairs, denoted as $\mathbf{E}_{tri}\in\mathbb{R}^{t\times 3\times d}$.
Figure 3: Illustration of our path tracking and aggregation mechanism.
@@
Utilizing the triplet pairs, $\mathbf{E}_{tri}$, as delineated in the preceding section, we introduce a novel path tracking and aggregation mechanism. This methodology is graphically showcased in Figure 3 and methodologically detailed in Algorithm 1. At its core, this mechanism is designed to cultivate a sophisticated interplay between the question and the dialog history. Notably, it is proficient in extracting pivotal semantic nuances from the dialog history, which subsequently enriches the representation of the question. This process is orchestrated through two consecutive phases: path tracking and path aggregation.
3.2.2 Path Tracking

### v2

# Uncovering Hidden Connections: Iterative Search and Reasoning for Video-grounded Dialog

@@ Abstract
In contrast to conventional visual question answering, video-grounded dialog necessitates a profound understanding of both dialog history and video content for accurate response generation. Despite commendable progress made by existing approaches, they still face the challenges of incrementally understanding complex dialog history and assimilating video information. In response to these challenges, we present an iterative search and reasoning framework, which consists of a textual encoder, a visual encoder, and a generator. Specifically, we devise a path search and aggregation strategy in the textual encoder, mining core cues from dialog history that are pivotal to understanding the posed questions. Concurrently, our visual encoder harnesses an iterative reasoning network to extract and emphasize critical visual markers from videos, enhancing the depth of visual comprehension. Finally, we utilize the pre-trained GPT-2 model as our answer generator to decode the mined hidden clues into coherent and contextualized answers. Extensive experiments on three public datasets demonstrate the effectiveness and generalizability of our proposed framework.
@@
Video-grounded Dialog, Path Search, Path Aggregation, Iterative Reasoning.
@@ 1 Introduction
With the rapid proliferation of multimedia data, research attention has focused on vision-language tasks, proving their pivotal role in academic research [1, 2, 3, 4, 5]. Notable advancements have been realized in fields such as Visual Question Answering (VQA) [6, 7, 8], Visual Dialog [9, 10], and Video Captioning [11, 12]. Central to these tasks is the aim to refine machine capabilities in understanding visual content and to convey these understandings linguistically in ways intelligible to humans.
@@
An emerging domain of significance is video-grounded dialog [^1] https://video-dialog.com/, sometimes referred to as audio-visual scene-aware dialog by certain academicians [13, 14, 15, 16]. As illustrated in Fig. 1, this rapidly evolving domain mandates the creation of an agent adept at conducting coherent dialogs rooted in video data, which is a monumental leap toward genuine human-computer synergy. The implications of video-grounded dialogue extend widely, enhancing virtual personal assistants and creating innovative solutions for visually impaired individuals, highlighting its significant academic relevance. Spurred by advancements in deep learning, the domain of video-grounded dialog is witnessing considerable traction. Existing techniques in this domain can be methodically categorized into:
@@
1) Textual Information Modeling. This branch emphasizes modeling textual elements, primarily involving question inquiries and their preceding contexts [17, 18, 19]. For instance, Le et al. [18] devised a semantic graph based on the lexical components of question-answer pairs, leading to a model proficient in forecasting reasoning pathways on the graph. This method is adept at extracting question-specific data from dialog contexts. Similarly, Kim et al. [17] employed the Gumbel-Softmax technique to identify dialog content that is most pertinent to a given question, enhancing question representation in the process.
@@
2) Visual Information Modeling. This kind of method highlights visual elements, primarily focusing on video frames and notable objects [20, 21, 22, 23]. As a case in point, Le et al. [21] propounded a bidirectional model that encapsulates both spatial-to-temporal relations and their converse, aiming to mine the complex spatial and temporal cues embedded in the video. Concurrently, Geng et al. [20] advocated representing videos via spatio-temporal scene graphs, capturing salient audio-visual cues and their semantic interrelations effectively.
@@
3) Decoder Generation. This section emphasizes the formulation of answer derivations [24, 25, 26]. For example, Le et al. [25] extended the established functionalities of the GPT-2 architecture, discarding conventional auto-regressive decoding strategy. Complementing this domain, the integration of pointer networks [24] has amplified the generative capabilities of dialog systems.
@@
Despite the significant progress observed in the domain of video-grounded dialog, there are still critical challenges that demand meticulous consideration: 1) Dialog History Understanding. Notable recent studies [19, 18] predominantly focus on dialog history segments that align directly with user queries. Such a perspective can inadvertently exclude broader dialog contexts, leading to the omission of significant information. The inherent lack of transparency in these methods further impedes a clear understanding of the decision-making process of the model. Hence, the imperative lies in developing a superior textual encoder capable of thoroughly analyzing dialog histories while improving model interpretability. 2) Visual Content Understanding. While several works [20] endeavor to extract visual signals through complex graph structures, the importance of user-posed questions in steering visual comprehension is frequently disregarded. This omission leads to a failure to capture question-specific insights, thereby limiting the capacity to mine meaningful visual information. Additionally, the current trend of depending solely on a singular extraction process [22, 23] may be inadequate considering the multi-faceted nature of video data. Therefore, an approach that harmonizes user queries with iterative reasoning is crucial for a nuanced comprehension of visual content.
@@
In response to the outlined challenges, we introduce the Iterative Search and Reasoning (ISR) framework for video-grounded dialog, detailed in Fig. 2. This novel approach aims to address the limitations of current methods by integrating an explicit history modeling strategy alongside an iterative reasoning network. To be specific, we implement a clear and structured path search and aggregation strategy, aimed at uncovering important information related to user queries within the dialogue context, thereby improving question comprehension. Concurrently, to harness the latent visual semantics in videos, a multimodal iterative reasoning network is devised. This architecture enables robust iterative reasoning processes, progressively refining the comprehension of visual content and consequently bolstering generation capabilities. The effectiveness of the ISR method is demonstrated through extensive experiments on three public datasets. To advance collaborative research, the codebase is available at https://github.com/Hyu-Zhang/ISR.
@@
  • We design an interpretable path search and aggregation strategy, which enables the effective extraction of crucial information from the preceding dialog context, thereby enhancing the comprehensive understanding of user intention.
  • We devise a multimodal iterative reasoning network that leverages user queries. This network serves to incrementally refine the comprehension of visual content, leading to more precise answer generation.
@@ 2.1 Visual Dialog
Visual dialog represents a sophisticated extension of the VQA framework, requiring multiple rounds of sequential interactions. The fundamental goal of this paradigm is to facilitate a seamless integration of cross-modal information derived from various inputs, culminating in the formation of a coherent textual response. Current methods within this domain may be broadly grouped into three distinct paradigms: attention mechanism-based strategies, graph network-based techniques, and approaches reliant on pre-trained models.
@@
Attention Mechanism-based Approaches. These approaches deploy attention mechanisms to pinpoint the visual regions that bear the highest relevance to the posed question. The underlying process requires an acute understanding of the immediate context of the inquiry, as well as an insight into the historical dialog trajectory. A salient example of this approach is the Recursive visual Attention (RvA) model, formulated by Niu et al. [27]. Through iterative refinement of visual attention and retrospective analysis of dialog sequences, this model adeptly addresses co-reference ambiguities. Further complementing this area, Yang et al. [28] have innovated two Sequential Dialog Networks (SeqDialN). These networks artfully separate the complexity inherent in multimodal feature fusion from the inference process, thereby simplifying the design and implementation of the inference engine.
@@
Graph Network-based Techniques. Recognizing the pivotal role of seamless information integration, the academic community is increasingly embracing graph network techniques [29]. For instance, Schwartz et al. [30] brought forth the Factor Graph Attention (FGA) architecture. This methodology enables the meticulous construction of graphs atop utility representations, subsequently discerning their interconnections. In a parallel vein, Guo et al. [31] introduced the Context-Aware Graph (CAG) neural structure, characterized by the dynamic interdependencies exhibited by individual nodes within the graph. Uniquely, this paradigm ensures that only nodes bearing the utmost relevance are integrated, yielding a refined, context-sensitive relational graph logic.
@@
Pre-trained Model-dependent Methods. The contemporary era of machine learning has been marked by the rise of pre-trained models, praised for their ability to seamlessly integrate diverse informational modalities. Anchoring on this insight, Wang et al. [32] delineated a holistic strategy that harnesses pre-trained architectures to adroitly weave together visual and dialogual elements. Mirroring this trajectory, Nguyen et al. [33] unveiled the Light-weight Transformer for Many Inputs (LTMI), a pioneering neural framework harnessing an agile transformer mechanism to masterfully orchestrate interactions across diverse utilities.
@@
However, the nuances and complexities of video-grounded dialog task preclude the direct application of methodologies cultivated for image-based visual dialog. Video-grounded dialog manifests longer utterance sentences, additional dialog rounds, and a more intricate context that demands sophisticated comprehension [34]. These attributes compound the challenge, rendering the direct transposition of existing visual dialog techniques impractical and inefficacious.
@@ 2.2 Video-grounded Dialog
In recent years, the academic community has observed an escalating focus on video-grounded dialog systems. Noteworthy challenges, including DSTC7 [34] and DSTC8 [35], have been instrumental in formulating benchmarks in this domain. This burgeoning field has subsequently witnessed the advent of numerous sophisticated methodologies to address its inherent complexities. For example, Hori et al. [13] pioneered an approach employing an LSTM-driven encoder-decoder framework, emphasizing multimodal attention to harmoniously integrate textual and visual cues. Expanding on this work, Le et al. [14] presented the Multimodal Transformer Network (MTN), devised specifically to encode visual content while concurrently integrating diverse data modalities. Similarly, innovative contributions like the Bi-directional Spatio-Temporal reasoning model (BiST)[21] have made strides in extracting both visual and spatial cues from videos, thus elucidating the interconnectedness of text and visual sequences in multidimensional aspects. In pursuit of refining co-referential understanding, Kim et al. [17] introduced the Structured Co-reference Graph Attention model (SCGA), laying the foundation for a graph-based approach deeply rooted in multimodal co-referencing techniques. In tandem with this, Geng et al. [20] advanced the Spatio-Temporal Scene Graph Representation learning technique (STSGR)—an innovative proposition capturing the semantic crux of videos through scene graphs, all the while maintaining efficient encoding.
@@
The ascendance of pre-trained language models in myriad natural language processing domains is unmistakable. Recent efforts are directed toward fusing these models with video-grounded dialog systems. To this end, Le et al. [25] advanced the VGD-GPT framework, leveraging the capabilities of the well-regarded GPT-2 model, thus transforming video-grounded dialog generation into a more fluid sequence-to-sequence task. Building on this progression, Li et al. [15] incorporated a multi-task learning approach with pre-existing language models, culminating in the inception of the Response Language Model (RLM). Recognizing the hurdles of text hallucination in text generation, Yoon et al. [26] crafted the Text HallucinAtion Mitigating (THAM) architecture, integrating Text Hallucination Regularization (THR) loss into expansive pre-trained language models. In a stride towards achieving nuanced dialog generation, Zhao et al. [36] advocated for a Multi-Agent Reinforcement Learning (MARL) methodology, accentuating sophisticated collaborations in cross-modal analysis. With the advent of Large Video-Language Models (LVLMs), such as Video-LLaMA [37] and Otter [38], it is evident that this intricate domain continues to be at the forefront of research attention.
@@
To sum up, within the domain of history modeling, a prevailing trend involves the deployment of intricate graph networks to discern semantic nuances intertwined within historical data. However, these methods often compromise on interpretability. In contrast, our approach embraces a streamlined path search and aggregation mechanism, offering superior interpretability. Additionally, current methodologies largely overlook the iterative interplay between videos and questions. We address this oversight, by devising an innovative iterative reasoning network that prioritizes this symbiotic relationship.
@@ 3.1 Problem Formulation
The objective of video-grounded dialog centers on generating a relevant answer $A_{t}$ through a thorough analysis of a video $V$, a caption $C$ crafted by a human to describe the video, its associated dialog history $H$, and current question $Q_{t}$. The dialog history $H$ comprises a sequence of prior questions and their answers, denoted as $(Q_{1},A_{1},Q_{2},A_{2},...,Q_{t-1},A_{t-1})$.
@@ 3.2.1 Preprocessing
The word embeddings for the given question $Q_{t}$, the caption $C$, and the dialog history $H$ are obtained through the same process. For clarify, we consider the question $Q_{t}=(w_{1},w_{2},...,w_{n_{q}})$ as an example. Here, an operation $f$ is applied to the vocabulary obtained from the word-to-id mapping function Vocab($\cdot$). This process is represented by the following expression:
@@
where $\mathbf{E}_{q}\in\mathbb{R}^{{n_{q}}\times{d}}$ is the resulting word embeddings of question $Q_{t}$. The $f$ corresponds to retrieving the embedding vector associated with the id from the pre-established embedding matrix. The same step is applied to get the feature of caption $C$ and dialog history $H$, symbolized as $\mathbf{E}_{c}\in\mathbb{R}^{n_{c}\times d}$ and $\mathbf{E}_{h}\in\mathbb{R}^{(t-1)\times n_{h}\times d}$, where $n_{c}$ and $n_{h}$ respectively denote the number of words present in the caption and each utterance [^2] A question-answer pair is considered a utterance..
@@
To improve the interpretability of questions drawn from the dialog history [^3] Given that the caption inherently encapsulates the essential information, it is utilized directly in the decoding generation process without necessitating further operations., we introduce two preprocessing steps: co-reference resolution and dependency parsing. Initially, the dialog history is integrated with the present question to create a unified sequence. We then utilize the AllenNLP [^4] https://allenai.org/allennlp. toolkit for co-reference resolution, aiming to dispel potential confusion related to pronouns. Following this, the Stanford CoreNLP package [^5] https://stanfordnlp.github.io/CoreNLP/. is employed to produce the dependency parse tree. Our attention is concentrated on the parsing of dialog responses due to the inherent similarities between questions and their corresponding answers. The dependency tree delineates the syntactic relationships between sentence constituents. By adopting the subject-verb-object paradigm, we extract pertinent triplets in the $\langle{subject,relation,object}\rangle$ format. This conversion facilitates the representation of both the question and dialog history as sets of triplet pairs, denoted as $\mathbf{E}_{tri}\in\mathbb{R}^{t\times 3\times d}$.
Figure 3: Illustration of our path search and aggregation strategy.
@@
Utilizing the triplet pairs $\mathbf{E}_{tri}$, we introduce a interpretable path search and aggregation strategy. This strategy is graphically showcased in Fig. 3 and methodologically detailed in Algorithm 1. It is designed to extract pivotal semantic information from the dialog history, enriching subsequently the representation of the question, which involves two consecutive phases: path search and path aggregation.
3.2.2 Path Search